What would be the solution and answer for this equation?
4x+3y=26
y=3x
y+5x=15
8y+5x=20
3y-9x=-18
y-3x=-3
x=4
22=y+2x

Answer by Alan3354(69443)

What would be the solution and answer for this equation?
4x+3y=26
y=3x
y+5x=15
8y+5x=20
3y-9x=-18
y-3x=-3
x=4
22=y+2x
----------
Is that one problem?